# High-Precision Simulations of the Parity Conserving Directed Percolation Universality Class in 1+1 Dimensions

### v1

@@ Abstract
Next to the directed percolation (DP) universality class, parity conserving directed percolation (pcDP; also called parity conserving branching annihilating random walks, pcBARW) is the second-most important model with an absorbing state transition. Its distinction from ordinary DP is that particle number is conserved modulo 2, which implies that there are two distinct sectors in systems with a finite initial number of particles: Realizations with even and odd particle numbers show different scaling behaviors, and systems in the odd sector cannot die. An intriguing feature of pcDP it is that some of its critical exponents seem to be very simple rational numbers. The most prominent is the one describing the average number of particles (or active sites) in the even sector, which is asymptotically constant. In contrast, the dynamical critical exponent (which is the same in both sectors) seems not close to any simple rational. Finally, the order parameter exponent $\beta$ (which is also the same in both sectors) is, according to the most precise previous simulations, rather close to 1, but incompatible with it. We present high statistics simulations which clarifies this situation, and which indicate several other intriguing properties of pcPD clusters. In particular, we find that all exponents which were close to rationals are even closer, and $\beta=1.000$ with the error in the next digit.
@@ I Introduction
The parity conserving directed percolation (pcDP) universality class [1, 2], also called even-offspring or parity conserving branching-annihilating random walks (pcBARW) [3, 4, 5, 6], directed Ising model [7, 8], generalized Domany-Kinzel model [9], or $\mathbb{Z}_{2}$ symmetric directed percolation (DP2) [10], was introduced in [1] as a counter example to the Janssen-Grassberger [11, 12] conjecture according to which absorbing state transitions with a unique absorbing state are always in the directed percolation (DP) universality class, provided they are continuous (‘2nd-order’) transitions. pcPD does not comply with that conjecture because it has two degenerate absorbing states. It is most interesting in one spatial dimension, where it allows for several physical interpretations.
The most straight forward is as directed percolation [13] (or, if time is continuous, the contact process [13]) with the added constraint that the particle number is conserved modulo 2 [1]. Due to this, any initial configuration with an odd number of particles on a lattice with periodic b.c. cannot die, which is sufficient to establish it as distinct from ordinary DP. The two absorbing (macro-)states in the sector with odd particle number are in this interpretation those where the leftmost active particle is at an even (odd) position. This immediately suggests another interpretation where the ‘particles’ of the contact process are domain walls [1]. If each lattice site is occupied by an Ising spin, then one can choose a suitable combination of spin flip and spin exchange (Kawasaki) moves so that the domain wall evolution mimics exactly the directed percolation [7]. Notice that this analogy between particle and domain wall models breaks down in higher dimensions, where particles still have dimension 0 while domain walls have dimension D1D-1.
@@
It is the main purpose of the present study to improve on this unsatisfactory situation by simulations with very high statistics, which might hopefully decide with reasonable safety whether the exponents in pcPD are simple or not, and find thereby any hint for a possible solvability of the model. For this purpose, we spent about 25 years of CPU time on work stations and laptops. We studied also several observables not analyzed in previous studies. We found some remarkable features and a possible hint why $\eta_{2}=0$, but no clear indication for solvability. On the other hand, we found clear evidence that $\beta=1$ exactly.
@@ II The model and computational details
In the last two simulations we did not find any surprises. Nor were the results sufficiently precise that we learned anything new. In order to keep the length of the paper reasonble, we will not present more details.
@@ III.2 Value of $p_{c}$ and Exponents Exactly at the Critical Point
For more precise statements we have to deal more carefully with the corrections to scaling, which are obviously not negligible. It is easy to see that the leading correction term has an exponent close to 1, but this term alone would not give an acceptable fit. In Fig. 2 we show a fit with three correction terms with exponents -1, -2, and -3. The amplitudes in front of these exponential terms are small and decrease with the rank, suggesting thus that the fit is indeed physically meaningful. Taking these scaling corrections into account, our best estimates for $p_{c}$ and $\eta_{2}$ are
$p_{c}=.450978(3),\quad\eta_{2}=0.0000(1).$ (14)
@@
An analogous analysis of $P_{2}(t)$ was much less successful. In Fig. 3 we show $t^{x}P_{2}(t)$, where $x=0.2872$ is the value of $\delta_{2}$ estimated in [14]. In contrast to $N_{2}(t)$, the curve $p=p_{c}$ has now no definite curvature. This means that definitely more than one correction term is needed for a quantitative fit. Also, the slow upward curvature for large $t$ suggests that the leading correction term has an exponent $\beta_{1}<1$ (roughly, we found $\beta_{1}\approx 0.5$), i.e. there are non-analytic corrections while all correction in $N_{2}(t)$ were compatible with being analytic. But the most troubling fact is that we were not able to produce acceptable fit, similar to Fig. 2.
@@
If $\eta_{2}=0$ as suggested by the data, then Eq.(7) says that $\delta_{2}=\eta_{1}$, so that we can estimate $\delta_{2}$ also from simulations starting from a single active site. In Fig. 4 we plot $t^{-\eta_{1}}N_{1}(t)$ against $t$, with three correction terms added. We see again a nearly perfect fit. It would suggest $p_{c}=.450982(5)$, which together with our previous estimate gives our best value $p_{c}=.450979(2)$. Taking into account this uncertainty of $p_{c}$, our best estimate of $\eta_{1}$ is
@@
In the fit shown in Fig. 4, the leading correction to scaling has an exponent -0.949. Taken at face value, this suggests a non-analytic correction term. But it also suggests that the leading term should actually have exponent -1 and is analytic. We were able to produce an acceptable (although not as good as in Fig. 4) fit with exponent -1, but only with four correction terms and with amplitudes which increase rapidly with rank, so that the amplitude for the term with exponent -4 is $|a_{4}|>10$. We thus conjecture that the leading correction is indeed non-analytic.
@@
The estimates for the dynamical exponent $z$ obtained from $R_{1}(t)$ and $R_{2}(t)$ were in perfect mutual agreement, and seemed to involve mostly analytic correction terms. We show results for $R^{2}_{1}(t)$ in Fig. 5, and for $R^{2}_{2}(t)$ in Fig. 6. As the previous plots these are log-linear plots. They show $R^{2}_{i}(t)/t^{2/z}$ (with $z$ being in each plot the best estimate) and with four correction terms, one of them being non-analytic with exponent -1.5 in both plots. The best estimate of $z$ combined from these two plots is
@@ III.6.2 The Order Parameter Exponent
Figure 18: (Color online) The highest six coefficients in the least square fits, for the same vales of $\beta$ as in the previous figure.
@@
Our obvious conjecture is that $\nu=2$ and $\beta=1$ exactly, and that $z=\nu_{t}/2$ is the only exponent which is not a very simple rational.
Our final estimates for the critical exponents are summarized, together with the values obtained in [18], in Table 1. We see that we could in all cases reduce the statistical error. Moreover, in all cases where the exponents are close to simple rationals, our estimates are closer to them that the estimates of [18]. Our final estimate of the critical point is $p_{c}=0.450976(3)$.
@@ IV Discussion and Conclusions
(iv) Different observables show vastly different corrections to scaling. While $N(t)$ and $R(t)$ in the even sector have corrections which are mostly or even entirely analytic (the leading correction exponent is 1 in both cases), others like $P_{2}(t)$ have exponents $\beta_{1}\approx 1/2$, and $Q_{2}^{\prime}(t)$ (the probability, in the even sector, to have a ‘pinch’ at time $t$) seems to have even $\beta_{1}=1/4$. Moreover, in these cases there are usually more than two significant correction terms needed for quantitative fits, which makes their determination rather difficult. We found no evidence that the exponents of the leading correction terms were different in the even and odd sectors, but found that they differed for different observables. This might suggest that the leading correction exponents are always the same, but coefficient of of the leading correction vanishes for some observables. A special role is played by $N(t)$ in the odd sector, where $\beta_{1}=0.949(5)$ which is close to $\beta_{1}=1$, but nevertheless clearly different.

### v2

@@ Abstract
Next to the directed percolation (DP) universality class, parity conserving directed percolation (pcDP; also called parity conserving branching annihilating random walks, pcBARW) is the second-most important model with an absorbing state transition. Its distinction from ordinary DP is that particle number is conserved modulo 2, which implies that there are two distinct sectors in systems with a finite initial number of particles: Realizations with even and odd particle numbers show different scaling behaviors, and systems in the odd sector cannot die. An intriguing feature of pcDP it is that some of its critical exponents seem to be very simple rational numbers. The most prominent is the one describing the average number of particles (or active sites) in the even sector, which is asymptotically constant. In contrast, the dynamical critical exponent (which is the same in both sectors) seems not close to any simple rational. Finally, the order parameter exponent $\beta$ (which is also the same in both sectors) is, according to the most precise previous simulations, rather close to 1, but incompatible with it. We present high statistics simulations which clarify this situation, and which indicate several other intriguing properties of pcPD clusters. In particular, we find that all exponents which were close to rationals are even closer, and $\beta=1.000$ with the error in the next digit.
@@ I Introduction
The parity conserving directed percolation (pcDP) universality class [1, 2], also called even-offspring or parity conserving branching-annihilating random walks (pcBARW) [3, 4, 5, 6], directed Ising model [7, 8], generalized Domany-Kinzel model [9], or $\mathbb{Z}_{2}$ symmetric directed percolation (DP2) [10], was introduced in [1] as a counter example to the Janssen-Grassberger [11, 12] conjecture according to which absorbing state transitions with a unique absorbing state are always in the directed percolation (DP) universality class, provided they are continuous (‘2nd-order’) transitions. In order to see why pcDP must be distinct from DP, it is most interesting in one spatial dimension, where it allows for several physical interpretations.
The most straight forward is as directed percolation [13] (or, if time is continuous, the contact process [13]) with the added constraint that the particle number is conserved modulo 2 [1]. Due to this, any initial configuration with an odd number of particles on a lattice with periodic b.c. cannot die, which is sufficient to establish it as distinct from ordinary DP. Conservation of particle numbers modulo 2 in 1 spatial dimension holds indeed in any model with two symmetric ground states, if the ‘particles’ are interpreted as domain walls. This was indeed the case for the two models studied in [1], which were stochastic cellular automata with $Z_{2}$ symmetries. The two absorbing states are in this realization periodic states $\ldots+-+$-$\dots$ and $\ldots-+-+\dots$, and the dynamics correspond to suitable combinations of spin flip and spin exchange (Kawasaki) moves [7]. Notice that this analogy between particle and domain wall models breaks down in higher dimensions, where particles still have dimension 0 while domain walls have dimension D1D-1.
@@
It is the main purpose of the present study to improve on this unsatisfactory situation by simulations with very high statistics, which might hopefully decide with reasonable safety whether the exponents in pcPD are simple or not, and find thereby any hint for a possible solvability of the model. For this purpose, we spent about 28 years of CPU time on work stations and laptops. We studied also several observables not analyzed in previous studies. We found some remarkable features and a possible hint why $\eta_{2}=0$, but no clear indication for solvability. On the other hand, we found clear evidence that $\beta=1$ exactly.
@@ II The model and computational details
In the last two simulations we did not find any surprises. Nor were the results sufficiently precise that we learned anything new. In order to keep the length of the paper reasonable, we will not present more details.
@@ III.2 Value of $p_{c}$ and Exponents Exactly at the Critical Point
For more precise statements we have to deal more carefully with the corrections to scaling, which are obviously not negligible. It is easy to see that the leading correction term has an exponent close to 1, but this term alone would not give an acceptable fit. In Fig. 2 we show a fit with three correction terms with exponents -1, -2, and -3. The amplitudes in front of these exponential terms are small and decrease with the rank, suggesting thus that the fit is indeed physically meaningful. Taking these scaling corrections into account, our estimates for $p_{c}$ and $\eta_{2}$ are
$p_{c}=.450979(3),\quad\eta_{2}=0.0000(1).$ (14)
@@
An analogous analysis of $P_{2}(t)$ was much less successful. In Fig. 3 we show $t^{x}P_{2}(t)$, where $x=0.2872$ is the value of $\delta_{2}$ estimated in [14]. In contrast to $N_{2}(t)$, the curve $p=p_{c}$ has now no definite curvature. This means that definitely more than one correction term is needed for a quantitative fit. Also, the slow upward curvature for large $t$ suggests that the leading correction term has an exponent $\beta_{1}<1$ (roughly, we found $\beta_{1}\approx 0.5$), i.e. there are non-analytic corrections while all correction in $N_{2}(t)$ were compatible with being analytic. But the most troubling fact is that we were not able to produce an acceptable fit, similar to Fig. 2.
@@
If $\eta_{2}=0$ as suggested by the data, then Eq.(7) says that $\delta_{2}=\eta_{1}$, so that we can estimate $\delta_{2}$ also from simulations starting from a single active site. In Fig. 4 we plot $t^{-\eta_{1}}N_{1}(t)$ against $t$, with three correction terms added. We see again a nearly perfect fit. It would suggest $p_{c}=.450981(5)$, which together with our previous estimate gives our preliminary best value $p_{c}=.450979(3)$. Taking into account this uncertainty of $p_{c}$, our best estimate of $\eta_{1}$ is
@@
In the fit shown in Fig. 4, the leading correction to scaling has an exponent -0.949. Taken at face value, this suggests a non-analytic correction term. But it also suggests that the leading term should actually have exponent -1 and is analytic. We were able to produce an acceptable (although not as good as in Fig. 4) fit with exponent -1, but only with four correction terms and with amplitudes which increase rapidly with rank, so that the amplitude for the term with exponent -4 is $|a_{4}|>10$. Thus the question is open whether the leading correction is non-analytic.
@@
The estimates for the dynamical exponent $z$ obtained from $R_{1}(t)$ and $R_{2}(t)$ are in perfect mutual agreement. The leading correction exponent is analytic with $\beta_{1}=1$. We show results for $R^{2}_{1}(t)$ in Fig. 5, and for $R^{2}_{2}(t)$ in Fig. 6. As the previous plots these are log-linear plots. They show $R^{2}_{i}(t)/t^{2/z}$ (with $z$ being in each plot the best estimate) and with four correction terms, one of them being non-analytic with exponent -1.5 in both plots. The best estimate of $z$ combined from these two plots is
@@ III.6.2 The Order Parameter Exponent
Figure 18: (Color online) The highest six coefficients in the least square fits, for the same values of $\beta$ as in the previous figure.
@@
Our obvious conjecture is that $\nu=2$ and $\beta=1$ exactly, and that $z=\nu_{t}/2=1/(2\delta_{2})$ is the only exponent which is not a very simple rational.
Our final estimates for the critical exponents are summarized, together with the values obtained in [18], in Table 1. We see that we could in all cases reduce the statistical error. Moreover, in all cases where the exponents are close to simple rationals, our estimates are closer to them that the estimates of [18]. Our final estimate of the critical point is $p_{c}=0.450977(2)$.
@@ IV Discussion and Conclusions
(iv) Different observables show vastly different corrections to scaling. While $N(t)$ and $R(t)$ in the even sector have corrections which are mostly or even entirely analytic (the leading correction exponent is 1 in both cases), others like $P_{2}(t)$ have exponents $\beta_{1}\approx 1/2$, and $Q_{2}^{\prime}(t)$ (the probability, in the even sector, to have a ‘pinch’ at time $t$) seems to have even $\beta_{1}=1/4$. Moreover, in these cases there are usually more than two significant correction terms needed for quantitative fits, which makes their determination rather difficult. We found no evidence that the exponents of the leading correction terms were different in the even and odd sectors, but found that they differed for different observables. This might suggest that the leading correction exponents are always the same, but coefficient of of the leading correction vanishes for some observables. An exception is $N(t)$ in the odd sector, where $\beta_{1}=0.949(5)$. This might suggest that the true exponent is $\beta_{1}=1$, but the question is open.